OpenStudy (anonymous):

Hello! How would you write this in logarithmic form? e^2 = 7.389

OpenStudy (lgbasallote):

Hint: \(\LARGE \log_{a} b = x \rightarrow a^x = b\)

OpenStudy (anonymous):

Yes, shouldn't it include In?

OpenStudy (unklerhaukus):

\[\ln x=\log_ex\]

OpenStudy (lgbasallote):

\(\Large \log_{e} a \rightarrow \ln a\)

OpenStudy (anonymous):

ln(7.389)=2 where ln is the natural logarithm (log base e)

OpenStudy (anonymous):

Thank you everyone! Have a good night!